Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by sud21 » Sat Jan 28, 2012 10:22 pm
A fair coin is tossed 5 times. What is the probability of getting at least three heads on consecutive tosses?
Join the discussion
Source: — Problem Solving |

by Mike@Magoosh » Sun Jan 29, 2012 3:20 pm
Hi, there. I'm happy to contribute to this. :)

This is a relatively challenging probability question.

First of all, toss a fair coin five times. Each toss has two possible outcomes, so the total number of five-toss outcomes is 2^5 = 32.

How many of those involve getting at least three heads on consecutive tosses?

Case 1: 3 H and 2 T

Here, the cases that satisfy the condition are:
HHHTT, THHHT, and TTHHH ----> 3 cases

Case 2: 4 H and 1 T

Here, the cases that satisfy the condition are:
HHHHT, HHHTH, HTHHH, HHHHT ----> 4 cases

Case 3: 5 H and 0 T

Here, there's only one case -- HHHHH -- and it satisfies the condition.

3 + 4 + 1 = 8 cases total.

P(at least three heads on consecutive tosses) = 8/32 = 1/4

Does that make sense? Please let me know if you have any questions.

Mike :)
Magoosh GMAT Instructor
https://gmat.magoosh.com/
Join the discussion

by Anurag@Gurome » Sun Jan 29, 2012 7:20 pm
sud21 wrote:A fair coin is tossed 5 times. What is the probability of getting at least three heads on consecutive tosses?
At least 3 heads implies it can 3 or 4 or 5 heads.

Case 1: Possibility of getting 3 heads on consecutive tosses
HHHTT, TTHHH, THHHT, HTHTH, HTHHH, HHHTH
So, the probability of getting 3 heads on consecutive tosses = 5 * (1/2)^5 = 5/32

Case 2: Possibility of getting 4 heads on consecutive tosses
HHHHT, THHHH
So, the probability of getting 4 heads on consecutive tosses = 2 * (1/2)^5 = 2/32

Case 3: Possibility of getting 5 heads on consecutive tosses
HHHHH
So, the probability of getting 5 heads on consecutive tosses = 1 * (1/2)^5 = 1/32

Therefore, required probability = 5/32 + 2/32 + 1/32 = 8/32 = [spoiler]1/4[/spoiler]

[spoiler][/spoiler]
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by gmatdriller » Mon Mar 17, 2014 2:54 pm
How does BINOMIAL EXPANSION work here?

(H + T)^5 = H^5 + 5H^4*T + 10H^3*T^2 + 10H^2*T^3 + 5H*T^4 + T^5

I feel we could collect just the terms that correspond to the number
of outcomes requested.

For example: An outcome of 4 Heads, 1 Tail => 5H^4*T
This is equal 5(1/2^5) = 5/32

But for an outcome of 3Heads, 2Tails = 10H^3*T^2 = 10/32 ..however, this is wrong
Correct answer = 5(favorable outome) / 32 (Total possible outome)

CAN SOMEONE PLEASE CORRECT WHERE I HAVE GONE WRONG
Join the discussion

by Mike@Magoosh » Mon Mar 17, 2014 3:23 pm
gmatdriller wrote:How does BINOMIAL EXPANSION work here?

(H + T)^5 = H^5 + 5H^4*T + 10H^3*T^2 + 10H^2*T^3 + 5H*T^4 + T^5

I feel we could collect just the terms that correspond to the number
of outcomes requested.

For example: An outcome of 4 Heads, 1 Tail => 5H^4*T
This is equal 5(1/2^5) = 5/32

But for an outcome of 3Heads, 2Tails = 10H^3*T^2 = 10/32 ..however, this is wrong
Correct answer = 5(favorable outome) / 32 (Total possible outome)

CAN SOMEONE PLEASE CORRECT WHERE I HAVE GONE WRONG
Dear gmatdriller,
I'm happy to respond. :-)

I believe you overlooked the crucial word "consecutive". If the question were just about 3+ heads in five tosses, then order would not matter at all, and the binomial expansion would be perfectly appropriate.

Instead, because we are concerned with 3+ consecutive heads, order means everything, and the binomial expansion is useless.

Does this make sense?
Mike :-)
Magoosh GMAT Instructor
https://gmat.magoosh.com/
Join the discussion

by gmatdriller » Wed Mar 19, 2014 9:25 pm
Hello Mike,
Here the Heads have to follow each other without interruption
in any Tails, right?

In that case, the example given by Sund21(shown below) would be wrong:
"Case 1: Possibility of getting 3 heads on consecutive tosses
HHHTT, TTHHH, THHHT, HTHTH, HTHHH, HHHTH"

I can see that if order matters, we would also have fewer no of cases.
In Binomial(order is not relevant)we have several pattern of 10.
In consecutive, there are only 5 ceases.

Please comment if my explanations are correct,

Thanks

Please
Join the discussion

by Mike@Magoosh » Thu Mar 20, 2014 9:41 am
gmatdriller wrote:Hello Mike,
Here the Heads have to follow each other without interruption
in any Tails, right?

In that case, the example given by Sund21(shown below) would be wrong:
"Case 1: Possibility of getting 3 heads on consecutive tosses
HHHTT, TTHHH, THHHT, HTHTH, HTHHH, HHHTH"

I can see that if order matters, we would also have fewer no of cases.
In Binomial(order is not relevant)we have several pattern of 10.
In consecutive, there are only 5 ceases.

Please comment if my explanations are correct,

Thanks

Please
Dear gmatdriller,
I think you misunderstand the exact meaning of the word "consecutive." The condition stated in the problem is "What is the probability of getting at least three heads on consecutive tosses?" In other words, three consecutive tosses, three tosses in a row, all have to be H. The string of five tosses must contain, somewhere, a streak of three H's in a row. As long as there are three H's a row, it doesn't matter if the others are head or tails and whether they are adjacent to the streak of three. For example, here are some scenarios that meet the condition:
HHHTT = YES
HHHTH = YES
THHHT = YES
TTHHH = YES
HTHHH = YES
Each one of those has a streak of three consecutive H's, so those correctly satisfy the condition. Of course, the condition allows for more than three consecutive heads, because of those important words "at least", so longer streaks are allowed:
HHHHT = YES
THHHH = YES
HHHHH = YES
Those eight cover all the possibilities in which a series of five coin tosses would have three or more consecutive H's. If the H's are split us, such that there are no streaks of three, then those don't meet the condition. For example
HTHTH = NO
HHTHH = NO
Both of those have 3 or more H's, but there is not a streak of three in a row, so they do not meet the condition.

Does all this make sense?

Mike :-)
Magoosh GMAT Instructor
https://gmat.magoosh.com/
Join the discussion

by gmatdriller » Mon Mar 24, 2014 1:58 pm
Thanks Mike for taking the time to explain further.

You seem, however, to explain just what I stated in my post.:

"HTHTH" is NOT a string of at least 3 consecutive Heads

HHHTT has 3 consecutive Heads with no interruption until after
the 3rd H.

In sum, Binomial expansion includes both the consecutives and
non-consecutives. And so, the Binomial option (10H3T^2) has more terms.

Thanks so much Mike.
Join the discussion